A bicyclist rode into the country for 5 h. In returning, her speed was 5 mi/h faster and the trip took 4 h. What was her speed eacy way?
Answer by ankor@dixie-net.com(22740)

A bicyclist rode into the country for 5 h. In returning, her speed was 5 mi/h faster and the trip took 4 h. What was her speed each way?
:
Let s = speed into the country
then
(s+5) = speed returning
:
The distance for both trips is the same. Write a distance equation
Dist = time * speed
:
Dist into the country = Dist back from the country
5s = 4(s+5)
:
5s = 4s + 20
:
5s - 4s = 20
:
s = 20 mph is into the country
:
20 + 5 = 25 mph back.
:
:
Check solution by finding if both trips are equal:
5*20 = 100
4*25 = 100
:
Make sense to you now?